# Non-linear Welfare-Aware Strategic Learning

### v1

@@ 6.3 Comparison with Previous Algorithms
6.3.2 Improve welfare does not ensure fairness.
The middle plot and the right plot of Fig. 4 compare the unfairness of different algorithms. Although STWF has the highest welfare among all algorithms, it does not ensure fairness. These results further shed light on a thought-provoking question: how likely is it for a fairness-aware algorithm to “harm" each party’s welfare under the strategic setting? Recent literature on fairness (Guldogan et al., 2022; Heidari et al., 2019) already began to consider “agent improvement" when strategic behaviors are present. However, it remains an intriguing direction to consider fairness and welfare together.
7 Conclusions and Societal Impacts
To facilitate socially responsible decision-making on strategic human agents in practice, our work is the first to consider the welfare of all parties in strategic learning under non-linear settings and the agents have specific information levels and respond to the decision policy by first estimating the decision policy and then best respond to the estimated policy. We formally studied the relationships between decision-maker welfare, social welfare, and agent welfare and revealed that it is non-trivial to balance the welfare of each party. Thus, we then proposed STWF, the first welfare-aware algorithm for general strategic learning problems.
Although our algorithm can effectively balance the welfare of different parties, it relies on the accurate estimation of agent’s response. In practice, the agent’s behavior can be highly complex. With incorrect estimations of agents’ behavior, our algorithm may fail and lead to unexpected outcomes. Thus, it is worthwhile for future work to collect real data from human feedback, and learn agent response function. Additionally, our experiments demonstrate that welfare and fairness may not imply each other, making it necessary to consider both notions theoretically to further promote trustworthy machine learning.
@@ Appendix A Proofs
$f(\mathbf{x^{\prime}})=$ $f(\mathbf{x})+\sum_{i=1}^{d}\frac{\partial f}{\partial x_{i}}(\mathbf{x})(x^{\prime}_{i}-x_{i})+\frac{1}{2!}\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}(\mathbf{x})(x^{\prime}_{i}-x_{i})(x^{\prime}_{j}-x_{j})+\cdots$ (3)
$+\frac{1}{K!}\sum_{i_{1}=1}^{d}\cdots\sum_{i_{K}=1}^{d}\frac{\partial^{K}f}{\partial x_{i_{1}}\cdots\partial x_{i_{K}}}(\mathbf{x})(x^{\prime}_{i_{1}}-x_{i_{1}})\cdots(x^{\prime}_{i_{K}}-x_{i_{K}})$
@@
1. Proof of the necessity: When (i) does not hold, $f^{*}$ is not equal to $h$ and the agents cannot perfectly estimate $f^{*}$ with $Q^{f^{*}}_{x}$. Therefore, if (ii) does not hold, this means there must exist a cost function $c(x,x^{\prime})$, under which we can find a region $A\subset\text{dom}(X)$ where $Q^{f^{*}}_{x}$ increases while $h$ decreases inside $A$. Thus, if we specify a cost function and an agent to satisfy the requirement that the agent can only change her features within $A$, then $f^{*}$ is unsafe, and we prove that either (i) or (ii) should at least hold.
@@ Appendix B Plots with error bars
(b) Comparisons under the ACSincome-CA dataset.
(c) Comparisons under the German Credit dataset.

### v2

@@ 6.3 Comparison with Previous Algorithms
Improve welfare does not ensure fairness.
The middle and the right plots of Fig. 4 compare the unfairness of different algorithms. Although STWF has the highest welfare, it does not ensure fairness. These results further shed light on a thought-provoking question: how likely is it for a fairness-aware algorithm to “harm" each party’s welfare under the strategic setting? Recent literature on fairness (Guldogan et al. 2022; Heidari, Nanda, and Gummadi 2019) already began to consider “agent improvement" when strategic behaviors are present. However, it remains an intriguing direction to consider fairness and welfare together.
7 Conclusion & Societal Impacts
To facilitate socially responsible decision-making on strategic human agents in practice, our work is the first to consider the welfare of all parties in strategic learning under non-linear settings when the agents have specific information levels. Our algorithm relies on the accurate estimation of agent response. In practice, the agent’s behavior can be highly complex. With incorrect estimations of agents’ behavior, our algorithm may fail and lead to unexpected outcomes. Thus, it is worthwhile for future work to collect real data from human feedback, and learn agent response function. Additionally, our experiments demonstrate that welfare and fairness may not imply each other, making it necessary to consider both notions theoretically to further promote trustworthy machine learning.
Acknowledgement
This material is based upon work supported by the U.S. National Science Foundation under award IIS-2202699, by grants from the Ohio State University’s Translational Data Analytics Institute and College of Engineering Strategic Research Initiative.
@@ Appendix A Proofs
$f(x^{\prime})=$ $f(x)+\sum_{i=1}^{d}\frac{\partial f}{\partial x_{i}}(x)(x^{\prime}_{i}-x_{i})+\frac{1}{2!}\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}(x)(x^{\prime}_{i}-x_{i})(x^{\prime}_{j}-x_{j})+\cdots$ (2)
@@
1. Proof of the necessity: When (i) does not hold, $f^{*}$ is not equal to $h$ and the agents cannot perfectly estimate $f^{*}$ with $Q^{f^{*}}_{x}$. Therefore, if (ii) does not hold, this means there must exist a cost function $c(x,x^{\prime})$, under which we can find a region $A\subset\text{dom}(X)$ where $Q^{f^{*}}_{x}$ increases while $h$ decreases inside $A$. Thus, if we specify a cost function and an agent to satisfy the requirement that the agent can only change her features within $A$ and when $P_{X}$ is any continuous distribution, then $f^{*}$ is unsafe, and we prove that either (i) or (ii) should at least hold.
@@ Appendix B Plots with error bars
(b) Comparisons under the German Credit dataset.
(c) Comparisons under the ACSincome-CA dataset.